OpenStudy (anonymous):

solve the equation for m. then find the value of m for each value of n, someone explain this, im confused 2m = -6n - 5 ; n = -2, 0, 2

OpenStudy (anonymous):

just put the value of n and then calculate value of m. like : for n= -2 2m= -6*(-2) - 5 m = 12 - 5 / 2 m = 7/2 like this you can put other value of n and calculate the value of m

OpenStudy (anonymous):

the question first ask you to solve therefore you must write it like this 2m=-6n-5 m=(-6n-5)/2 Then write when n=-2: m=(-6 *(-2)-5)/2 =7/2 continue to do the same but with the other values for n :D